# Effect of Dark matter and $\sigma$-cut potential on radial and non-radial oscillation modes in neutron stars

### v1

@@ I Introduction
Detection of oscillation spectra provides key information on NS mass, frequency, tidal Love numbers, damping times, and moments of inertia, helping to constrain the EoS of dense matter [72]. Newly formed NSs in supernovae oscillate at characteristic frequencies determined by the restoring force, which may have different origins [73, 74, 75]. These oscillations are classified as radial and non-radial modes. While radial modes do not directly produce gravitational waves (GWs), their coupling with non-radial modes enhances GW emission, improving detectability [76, 77]. Chirenti et al. [78] noted that BNS post-merger remnants can form hyper-massive NSs, influencing gamma-ray bursts and radial oscillations. Their high-frequency modes (1similar-toabsent1$\sim$ 1∼ 14444 kHz) may be detectable by LIGO, Virgo, and KAGRA [79, 80], with improved sensitivity in third-generation detectors [68, 69, 70].
@@ III.1 Non-radial
×$\times\Bigg{\{}(1-e^{\lambda}\textrm{Q})H_{0}+(r^{2}\omega^{2}e^{-\nu}+n_{l}+1)H_{1}$
@@ IV.2 Non-radial profiles
which can be seen in Fig. 4 by the brown dash-dotted line (Our Fit). The lower panel displays the relative error (%) between the calculated $f$-mode frequencies from the data and the best-fit line proposed in this work. Each marker corresponds to a NS configuration based on a given EoS. Solid markers represent soft EoS, while triangular markers denote stiff EoSs. The plot shows that the relative error remains below approximately 2.5% throughout the range of mean stellar densities considered. This indicates a strong agreement between the fit and the numerical data across all three EoS models analyzed. The clustering of most points below the 2% threshold supports the quasi-universal character of the $f$$\sqrt{M/R^{3}}$ relation, even in the presence of dark matter and $\sigma$-cut potential modifications to the EoS.
@@ IV.4 Radial profiles
Figure 7: Left: The radial displacement perturbation ${\xi(r)}$ = ${\Delta r/r}$ as a function of dimensionless radius distance ${r/R}$. The colorbar represents the modes: ${f}$-mode ($n=0$) and lower-order ${p}$-modes ($n=1-5$) for soft NL, NL-$\sigma$ cut, and NL DM EoS. Right: Same as left plot, but for stiff EoSs. All the modes are calculated at 1.4 ${M_{\odot}}$.
Figure 8: Left: The radial pressure perturbation ${\eta(r)}$ = ${\Delta r/r}$ as a function of dimensionless radius distance ${r/R}$ The colorbar represents the modes: ${f}$-mode ($n=0$) and lower-order ${p}$-modes ($n=1-5$) for soft NL, NL-$\sigma$ cut, and NL DM EoS. Right: The same as the left plot, but for stiff EoSs. All modes are calculated at 1.4 ${M_{\odot}}$.
Solving the Sturm-Liouville eigenvalue equations for $\omega$, we calculate the first 10 modes $\omega_{n}$, with $n=0$ corresponding to the $f$-mode and $n=1-9$ as high-order $p$-modes. The radial displacement perturbation profile $\xi(r)$ as a function of the dimensionless radial coordinate $r/R$ is shown in Fig. 7. The left (right) panel presents the profile for soft (stiff) NL, NL-$\sigma$ cut, and NL DM EoS. We display only the fundamental $f$-mode ($n=0$) and the first few lower-order $p$-modes ($n=1-5$) in the plot, with the colorbar representing the modes. The solid, dashed, and dot-dashed lines correspond to the NL, NL-$\sigma$ cut, and NL DM EoS, respectively. In the region $0<r<R$, the $n$th mode exhibits exactly $n$ nodes in the $\xi(r)$ profile, following the characteristics of a Sturm-Liouville system. The amplitude of $\xi_{n}(r)$ for each frequency mode $\nu_{n}$ is largest near the center and diminishes toward the surface. Lower-order modes exhibit a smooth decline, whereas higher-order modes display small oscillations, which become more pronounced for even higher modes.
@@
Fig. 9 illustrates the frequency separation, $\Delta\nu$ (in kHz), as a function of the oscillation frequency, $\nu$ (in kHz), for NS matter modeled by three different EoSs: NL, NL-$\sigma$ cut, and NL DM. The frequencies for the fundamental mode ($n=0$) and the higher $p$-modes ($n=1-9$) are displayed in Table 4. In this figure, the solid lines represent softer EoSs, while the dashed lines correspond to stiffer ones. The variations observed in $\Delta\nu$ as $\nu$ increases can be primarily attributed to the transition between the crust and the core of the neutron star, as discussed in Ref. [83, 84, 86, 89]. In the crust, where the density is relatively low and nuclei coexist with free neutrons, the oscillation modes are influenced by the inhomogeneous composition. As the density exceeds a critical threshold, the matter becomes uniform in the core, and this crust–core boundary gives rise to mode coupling or conversion effects that result in fluctuations in the oscillation frequencies. Softer EoSs, which produce steeper density gradients and more compact stars, tend to enhance the interaction between crustal and global modes, leading to more pronounced fluctuations in $\Delta\nu$. Conversely, stiffer EoSs generate smoother transitions in density and support larger masses and radii, shifting the oscillation modes and typically shortening or decreasing the frequency separation.
@@
Figure 9: Frequency difference ${\Delta\nu_{n}}$ = ${\nu_{n+1}}$ ${\nu_{n}}$ versus ${\nu_{n}}$ in kHz for NL, NL-$\sigma$ cut, and NL DM EoS. The solid (dashed) lines correspond to the soft (stiff) EoSs.

### v2

@@ I Introduction
Detection of oscillation spectra provides key information on NS mass, frequency, tidal Love numbers, damping times, and moments of inertia, helping to constrain the EoS of dense matter [72]. Newly formed NSs in supernovae oscillate at characteristic frequencies determined by the restoring force, which may have different origins [73, 74, 75]. These oscillations are classified as radial and non-radial modes. While radial modes do not directly produce gravitational waves (GWs), their coupling with non-radial modes enhances GW emission, improving detectability [76, 77]. Chirenti et al. [78] noted that BNS post-merger remnants can form hyper-massive NSs, influencing gamma-ray bursts and radial oscillations. Their high-frequency modes ($\sim 1$44 kHz) may be detectable by LIGO, Virgo, and KAGRA [79, 80], with improved sensitivity in third-generation detectors [68, 69, 70].
@@ III.1 Non-radial
$\times\Bigg{\{}(1-e^{\lambda}\textrm{Q})H_{0}+(r^{2}\omega^{2}e^{-\nu}+n_{l}+1)H_{1}$
@@ IV.2 Non-radial profiles
[…]
@@ IV.4 Radial profiles
Figure 7: Left: The radial displacement perturbation $\xi(r)$} = $\Delta r/r$} as a function of dimensionless radius distance r/R{r/R}. The colorbar represents the modes: f{f}-mode ($n=0$) and lower-order p{p}-modes (n=15n=1-5) for soft NL, NL-$\sigma$ cut, and NL DM EoS. Right: Same as left plot, but for stiff EoSs. All the modes are calculated at 1.4 $M_{\odot}$}.
Figure 8: Left: The radial pressure perturbation $\eta(r)$} = $\Delta r/r$} as a function of dimensionless radius distance r/R{r/R} The colorbar represents the modes: f{f}-mode ($n=0$) and lower-order p{p}-modes (n=15n=1-5) for soft NL, NL-$\sigma$ cut, and NL DM EoS. Right: The same as the left plot, but for stiff EoSs. All modes are calculated at 1.4 $M_{\odot}$}.
Solving the Sturm-Liouville eigenvalue equations for $\omega$, we calculate the first 10 modes $\omega_{n}$, with $n=0$ corresponding to the $f$-mode and n=19n=1-9 as high-order $p$-modes. The radial displacement perturbation profile $\xi(r)$ as a function of the dimensionless radial coordinate $r/R$ is shown in Fig. 7. The left (right) panel presents the profile for soft (stiff) NL, NL-$\sigma$ cut, and NL DM EoS. We display only the fundamental $f$-mode ($n=0$) and the first few lower-order $p$-modes (n=15n=1-5) in the plot, with the colorbar representing the modes. The solid, dashed, and dot-dashed lines correspond to the NL, NL-$\sigma$ cut, and NL DM EoS, respectively. In the region $0<r<R$, the $n$th mode exhibits exactly $n$ nodes in the $\xi(r)$ profile, following the characteristics of a Sturm-Liouville system. The amplitude of $\xi_{n}(r)$ for each frequency mode $\nu_{n}$ is largest near the center and diminishes toward the surface. Lower-order modes exhibit a smooth decline, whereas higher-order modes display small oscillations, which become more pronounced for even higher modes.
@@
Fig. 9 illustrates the frequency separation, $\Delta\nu$ (in kHz), as a function of the oscillation frequency, $\nu$ (in kHz), for NS matter modeled by three different EoSs: NL, NL-$\sigma$ cut, and NL DM. The frequencies for the fundamental mode ($n=0$) and the higher $p$-modes (n=19n=1-9) are displayed in Table 4. In this figure, the solid lines represent softer EoSs, while the dashed lines correspond to stiffer ones. The variations observed in $\Delta\nu$ as $\nu$ increases can be primarily attributed to the transition between the crust and the core of the neutron star, as discussed in Ref. [83, 84, 86, 89]. In the crust, where the density is relatively low and nuclei coexist with free neutrons, the oscillation modes are influenced by the inhomogeneous composition. As the density exceeds a critical threshold, the matter becomes uniform in the core, and this crust–core boundary gives rise to mode coupling or conversion effects that result in fluctuations in the oscillation frequencies. Softer EoSs, which produce steeper density gradients and more compact stars, tend to enhance the interaction between crustal and global modes, leading to more pronounced fluctuations in $\Delta\nu$. Conversely, stiffer EoSs generate smoother transitions in density and support larger masses and radii, shifting the oscillation modes and typically shortening or decreasing the frequency separation.
@@
Figure 9: Frequency difference $\Delta\nu_{n}$} = $\nu_{n+1}$} - $\nu_{n}$} versus $\nu_{n}$} in kHz for NL, NL-$\sigma$ cut, and NL DM EoS. The solid (dashed) lines correspond to the soft (stiff) EoSs.